The measure of the largest angle in a triangle is forty degrees larger than the smallest angle. The measure of the middle angle is twice as large as the smallest angle. Find the measurement of the three angles in the triangle.
Answer by TimothyLamb(4379)

x = largest angle (degrees)
y = middle angle (degrees)
z = smallest angle (degrees)
---
x = z + 40
y = 2z
x + y + z = 180
---
put the system of linear equations into standard form
---
x - z = 40
y - 2z = 0
x + y + z = 180
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 75
y = 70
z = 35
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---